Linearize y = aexp(-bx2)
and determine the coefficients a and
b, and the coefficient of determination using the data below.
Estimate the value of y at x= 5.5.

x 1 2 3 4 5 6 7 8 9
y 4.852 4.299 3.571 2.713 1.904 1.367 0.895 0.430 0.365
Answer #1

I have explained each and every step in the solution even the very basic ones.

After solving for a while, I came to know that I have done some extra steps that's why I have cut them in the solution. So no need to worry.

Please rate if you like !!!

This helps to keep us motivated :)

Ask doubts (if any) in the comments section.

9 = ae-by- thy = lna - bx2 let lay=Y It will solre using method of EY=anon -bex² Exy= ena Ex-bEx3 RY - least squares - - 14.8